Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx+b.
(3,-6); 5x+2y=9
Help me please.

Answer by scott8148(6628)

subtracting 5x __ 2y=-5x+9 __ dividing by 2 __ y=(-5/2)x+(9/2)

prependicular lines have slopes that are negative reciprocals
__ the slope of the given line is -5/2, so the slope of the new line is 2/5

using point-slope __ y-(-6)=(2/5)(x-3) __ y=(2/5)x-(36/5)